HELP!!!
If there were no air resistance, then the height
(in feet) above the earth for a skydiver t seconds after
jumping from an airplane at 10,000 feet would be given by
h(t)16t2  10,000.
a) Find the time that it would take to fall to earth with no
air resistance; that is, find t for which h(t)  0. A
skydiver actually gets about twice as much free fall
time due to air resistance.
b) Use the accompanying graph to determine whether the
skydiver (with no air resistance) falls farther in the first
5 seconds or the last 5 seconds of the fall.
c) Is the skydiver’s velocity increasing or decreasing as
she falls?

Answer by stanbon(75887)

If there were no air resistance, then the height
(in feet) above the earth for a skydiver t seconds after
jumping from an airplane at 10,000 feet would be given by
h(t) = 16t2 + 10,000.
------------------------------
a) Find the time that it would take to fall to earth with no
air resistance; that is, find t for which h(t) = 0.
---
-16t^2+10000 = 0
16t^2 = 10000
t = sqrt(10000/16)
t = 100/4
t = 25 seconds.
-------------------------

A skydiver actually gets about twice as much free fall
time due to air resistance.
b) Use the accompanying graph to determine whether the
skydiver (with no air resistance) falls farther in the first
5 seconds or the last 5 seconds of the fall.
---
No graph shown.
Find h(5 seconds) and h(20 seconds) and you should be able
to answer the question.
---------------------
c) Is the skydiver’s velocity increasing or decreasing as
she falls?
Look at the distance covered in the 5 sec. compared to the
distance covered in the last 5 sec.
=====================================
Cheers,
Stan H.
=====================================